square root of x+5+2=square root of 6x+2

Answer by edjones(8007)

sqrt(x+5+2)=sqrt(6x+2)
x+5+2=6x+2 Square each side.
x+7=6x+2
x-x+7-2=6x-x+2-2
5=5x
5x/5=5/5
x=1
.
Check:
sqrt(1+5+2)=sqrt(6+2) true
.
Ed